Mr. Ahmed retired as president of the Master Tile Company but is currently on consulting contract for Rs. 450,000 per year for next 10 years. a. If Mr. Ahmed’s opportunity cost is 10 percent, what is the present value of his consulting contract? b. Assuming that Mr. Ahmed will not retire for two more years and will not start to receive his 10 payments until the end of third year, what would be the value of his deferred annuity? 

 

 


1
Expert's answer
2021-02-24T10:06:40-0500

a)

"PV_{A}=A\times PV_{IFA}"

"(10\%, 10 periods)"

"=\$45000\times6.145"

"=\$276,525"


b)

"PV=FV\times PV_{IF}"

"\$276,525\times 0.826"

"=\$228,410"